OpenStudy (anonymous):

If y= -6e^(-0.25x)*cos(5.2x), help me solve the equation y'=0? :D

OpenStudy (anonymous):

y'=-6e^(-0.25x)(-0.25)*cos(5.2x) + (-6e^(-0.25x))(-sin(5.2x))(5.2) = = 1.5*e^(-0.25x)*cos(5.2x) + 31.2*e^(-0.25x)*sin(5.2x)

OpenStudy (anonymous):

= e^(-0.25x)((1.5*cos(5.2x) + 31.2*sin(5.2x))

OpenStudy (anonymous):

Don't know if I'm on the right track even :l